OpenStudy (anonymous):

I need help finding the proportion of the variable. y/12=77/84 I got 6.4166666 on my calculator and I know its not right.

OpenStudy (anonymous):

y= (12)(77)/(84)= 11 one third/3= 1/9. thus t= 6/(1/9)= 6x9=54

OpenStudy (anonymous):

Thank you that really helped alot.

OpenStudy (anonymous):

welcome